OpenStudy (anonymous):

find the area of a polygon with vertices at (1,3), (7,3), (7,7), and (4,7).

OpenStudy (anonymous):

Two ways to do it. If you plot the points, it shows a trapezium with parallel sides 6 (=7-1) and 3(=7-4) units long, and the height is 4 (=7-3). So the area is (6+3)*4/2=18.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yup! u could do it this way 2:If you would like to calculate the area of a polygon given in corner coordinates, you can list the points in sequence like a,b c,d e,f g,h a,b and repeat a,b at the bottom. Calculate the crossed sum for each successive pair of coordinates, as in: ad-bc + cf-de + eh-fg + gb-ha and divide the sum by 2. The sum may be negative depending the direction of the points. So just take the absolute value. For the given points, list: 1,3 7,3 7,7 4,7 1,3 (repeat of first point) Products: 1*3-3*7 + 7*7-3*7 + 7*7-7*4 + 4*3-7*1 =-18 + 28 + 21 + 5 = 36 Divide by 2 to get 18.

OpenStudy (anonymous):

medal me?;)

OpenStudy (anonymous):

How about find the perimeter of the polygon with vertices at (1,3), (7,3), (7,7), and (4,7).

OpenStudy (anonymous):

hold on..

OpenStudy (anonymous):

done?

OpenStudy (anonymous):

perimeter...(1,3) ---> (7,3) You remained at height y = 3, and went from x = 1 to x = 7, so horizontally you went 6 to the right. (7,3)----->(7,7) x = 7 in the first and x = 7 in the second, you didn't walk anything in the x direction, your height y changed from 3 to 7, so you went vertically up 4. (7,7) --> (4,7) You height remained the same, you went from 7 to 4 in the x direction, that's 3 to the left. Finally, to get back to point (1,3) - (4,7) ----> (1,3), 3 to the left and 4 down, the diagonal being 5. 6+4+3+5= 10 + 8 = 18... get it?

OpenStudy (anonymous):

done;)

OpenStudy (anonymous):

The perimeter of any polygon is the sum of the lengths of all the sides.

OpenStudy (anonymous):

medal me?;)

OpenStudy (anonymous):

ok and i need two more hold up

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

got to go.. sry!

OpenStudy (anonymous):

send me the rest and i'll try to get 2 them:)